OpenStudy (anonymous):

y=2x+6 2x-y=2 solve please?

jimthompson5910 (jim_thompson5910):

2x-y=2 2x-(2x+6)=2 2x-2x-6=2 -6=2 Since the last equation is FALSE, this means that there are no solutions.

OpenStudy (anonymous):

x = 2/7 y = 2